# Auditing Large Language Models for Enhanced Text-Based Stereotype Detection and Probing-Based Bias Evaluation

### v1

@@ Abstract
Recent advancements in Large Language Models (LLMs) have significantly increased their presence in human-facing Artificial Intelligence (AI) applications. However, LLMs could reproduce and even exacerbate stereotypical outputs from training data. This work introduces the Multi-Grain Stereotype (MGS) dataset, encompassing 51,867 instances across gender, race, profession, religion, and stereotypical text, collected by fusing multiple previously publicly available stereotype detection datasets. We explore different machine learning approaches aimed at establishing baselines for stereotype detection, and fine-tune several language models of various architectures and model sizes, presenting in this work a series of stereotypes classifier models for English text trained on MGS. To understand whether our stereotype detectors capture relevant features (aligning with human common sense) we utilise a variety of explanainable AI tools, including SHAP, LIME, and BertViz, and analyse a series of example cases discussing the results. Finally, we develop a series of stereotype elicitation prompts and evaluate the presence of stereotypes in text generation tasks with popular LLMs, using one of our best performing previously presented stereotypes detectors. Our experiments yielded several key findings: i) Training stereotype detectors in a multi-dimension setting yields better results than training multiple single-dimension classifiers.ii) The integrated MGS Dataset enhances both the in-dataset and cross-dataset generalisation ability of stereotype detectors compared to using the datasets separately. iii) There is a reduction in stereotypes in the content generated by GPT Family LLMs with newer versions.  [^1] The code, dataset, and model for this paper will be publicly released upon acceptance.
@@ 1 Introduction
The field of Artificial Intelligence (AI) continues to evolve with Large Language Models (LLMs) showing both potential and pitfalls. This research explores the ethical dimensions of LLM auditing in Natural Language Processing (NLP), with a focus on text-based stereotype classification and bias benchmarking in LLMs. The advent of state-of-the-art LLMs including OpenAI’s GPT series (Brown et al., 2020; Radford et al., 2019; OpenAI, 2023), Meta’s LLaMA series (Touvron et al., 2023a; b), and the Falcon series (Almazrouei et al., 2023) has magnified the societal implications. These LLMs, shown up with abilities like in-context learning as a few-shot learner (Brown et al., 2020), reveal emergent capabilities with increasing parameter and training token sizes (Wei et al., 2022). However, they show fairness concerns due to their training on extensive, unfiltered datasets such as book (Zhu et al., 2015) and Wikipedia corpora (Foundation, 2021b), and large internet corpora like Common Crawl (Foundation, 2021a). This training data often exhibits systemic biases and could further lead to detrimental real-world effects, confirmed by studies (May et al., 2019b; Bordia & Bowman, 2019; Davidson et al., 2019; Magee et al., 2021). For instance, biases in LLMs and AI systems can reinforce political polarization as seen in Meta’s news feed algorithm (Bakshy et al., 2015), and exacerbate racial bias in legal systems as documented in predictive policing recidivism algorithms like COMPAS (Angwin et al., 2016). Furthermore, issues such as gender stereotyping and cultural insensitivity are highlighted by tools like Google Translate and Microsoft’s Tay (Prates et al., 2019; Neff & Nagy, 2016). Most existing studies focus on either bias benchmarks in LLMs or text-based stereotypes detection and overlook the interaction between them, which remains underexplored and indicates gaps. Our study makes a clear line between Bias, as observable deviations from neutrality in LLM downstream tasks, and Stereotype, a subset of bias entailing generalized assumptions about certain groups in LLM outputs. Aligning with established stereotype benchmark: StereoSet (Nadeem et al., 2020), we detect text-based stereotypes at sentence granularity, across four societal dimensions—Race, Profession, Religion, and Gender—within text generation task conducted with LLMs.
@@ 2 Related Works
Further expanding on the benchmark, StereoSet (Nadeem et al., 2020) and CrowS-Pairs (Nangia et al., 2020) are popular dataset-based stereotype benchmarking approaches that use the examples in the datasets to calculate the masked token probabilities and pseudo-likelihood-based scoring of the LLM to assess whether stereotypical results are output. A key disadvantage of these approaches is that the stereotype assessment’s generalisation bounds are limited to the diversity of the examples in the datasets. On the contrary, we use these examples to fine-tuned models to detect stereotypes from any generated text. This gives our approach the advantage of assessing the LLM’s bias based on any text output generated by the LLM rather than within the constraints of the labelled datasets. Benchmarks such as WinoQueer (Felkner et al., 2023) and SeeGULL (Jha et al., 2023) focus on stereotype types that are out of the scope of this work (e.g. regional stereotype etc.). Benchmarks such as WEAT (Caliskan et al., 2017) and SEAT (May et al., 2019a) use pre-defined attribute and target word sets to assess stereotypical language, making them similar to StereoSet (Nadeem et al., 2020), WinoQueer (Felkner et al., 2023), and CrowS-Pairs (Nangia et al., 2020) approaches exposed to the same limitations. Meanwhile, the BBQ benchmark (Parrish et al., 2022), which centers on a question-answering task, has been observed to prompt LLMs to avoid answering due to its explicit question design. BOLD (Dhamala et al., 2021), focusing on text generation, employs various metrics but lacks a solid measure to identify stereotypes. These prior collectively reveal a significant gap in current methods, as they do not facilitate a high granular stereotype detection in freely generated text by LLMs.
Additionally, several other prior works (Bolukbasi et al., 2016; May et al., 2019b) could be used to implement token-level stereotype detection that is out of scope for this work as we focus on sentence-level stereotype detection. Albeit, these works also lack transparency, a gap our work addresses through eXlainable AI (XAI) techniques. While emerging LLM evaluation frameworks like DeepEval (Team, 2023), HELM (Liang et al., 2022), and LangKit (whylabs, 2023) takes a holistic view on bias evaluation, our framework complements them as our proposal can become a subcomponent within their systems.
@@ 3 Methodology
  • RQ3: How does our multi-class classifier built for stereotype detection compare to other relevant baselines?
@@ 3.1 MGS Dataset (RQ1-3)
The Multi-Grain Stereotype Dataset (MGSD) was constructed from two crowdsourced sources, StereoSet Nadeem et al. (2020) and CrowS-Pairs Nangia et al. (2020). It consists of 51,867 instances, divided into training and testing sets in an 80:20 ratio. This division ensures stratified sampling across different stereotype types and data sources, providing a diverse range of examples for model creation and facilitating multi-class learning. The dataset includes columns for original text (text), text with marked stereotypical words/phrases (text_with_marker), stereotype type (stereotype_type), stereotype level (category), original data source (data_source), and gold label (label).
The labelling scheme in the MGS Dataset classifies texts into three categories: ”stereotype”, ”neutral”, and ”unrelated”, covering four social stereotype dimensions: ”race”, ”religion”, ”profession”, and ”gender”. There are nine label types: ”unrelated”, ”stereotype_race”, ”stereotype_gender”, ”stereotype_profession”, ”stereotype_religion”, ”neutral_race”, ”neutral_gender”, ”neutral_profession”,”neutral_religion”. The dataset supports both sentence-level and token-level classification tasks. For preprocessing, the text was tokenized, and ”$=$” markers were inserted to highlight stereotypical tokens (e.g., He is a doctor $\rightarrow$ He is a $=$doctor$=$). This approach allows for the future use of the dataset in training token-level stereotype detectors and aids in generating prompts and counterfactual scenarios for sentence-level detector model evaluations.
StereoSet includes two types of data examples: (i) intra-sentence, where bias is within a single sentence, and (ii) inter-sentence, where bias spreads across multiple sentences. (iii) CrowS-Pairs features pairs of sentences that carry the stereotype or stereotype with counterfactual identity/group. In (i) intra-sentence cases, the correlated label is assigned to the single sentence. In (ii) inter-sentence cases, sentences are merged, and the label is assigned to create the final MGS dataset. For (iii) CrowS-Pairs, the first pair of sentences and the ”sent_more” feature are chosen, and the label is assigned to finalize the dataset. The MGS Dataset serves as a valuable resource for evaluating and understanding stereotypes. We conducted exploratory data analysis and provided details in Appendix A.4.
@@ 3.2 Finetuning the Stereotype Classifier and Explaining It (RQ1-4)
For RQ3, in order to compare the new model with comparative baselines, we developed popular machine learning methods as baselines, in addition to the range of PLMs we had fine-tuned. This was necessary as we could not find suitable multi-class baselines in previous research. We implemented the i) Random selection, that assigns labels at random, ii) a Logistic regression, and iii) Kernel SVM (sigmoid kernel identified empirically) models trained TF-IDF features from MGS Dataset. iiii) Finally, zero-shot classification task, we choose the ”sileod/deberta-v3-base-tasksource-nli” DeBERTa-based model that has shown the best performance in this task(Sileo, 2023).
For RQ4, to ensure robust validation and interpretation of our stereotype classifier, we employ multiple explainability ai methods for feature attribution and model structural interpretability. This allows us to check for consistency of explanations as different explainability methods can yield varying results in feature importance (Swamy et al., 2022). Specifically, we apply SHAP (Lundberg & Lee, 2017) and LIME (Tulio et al., 2016), two popular model-agnostic explainability techniques, to identify the text tokens most influential in the classification process. We use randomly selected examples from the test set of MGS Dataset to analyse explanations. Additionally, we utilize BERTViz (Vig, 2019), a model-specific visualization tool for transformer models, to observe how the model’s attention heads engages with specific tokens across layers.
@@ 3.3 Stereotype Elicitation Experiment and Bias Benchmarks (RQ5)
We first developed a method for prompt generation, resulting in a prompt library in Appendix A.3, which effectively elicits stereotypical text from the LLMs. We take examples from the MGS dataset and use the markers to identify the prompts (the part of the example before the marker) for the LLM under investigation. When selecting examples for generating prompts, we use word count-based prioritization logic, where initially, we target long examples resulting in detailed prompts. We generate prompts from the dataset for the different societal dimensions ($\approx 200$ per dimension). We further validate the neutrality of the identified prompts using the proposed model to ensure that all prompts have been classified as ”unrelated”. Then, we use the prompts library to probe the LLM under investigation to complete the rest of the passage (prompt). To evaluate the effectiveness of our prompts in eliciting stereotype-related content, we conducted perplexity tests, as documented in Appendix A.2. These tests compare the LLMs’ responses against a stratified sample from the MGS Dataset, based on labels. Finally, we use the generated output to detect stereotypes, which is the final assessment.
In this paper, we assess the bias of the GPT series of LLMs, considering only stereotype and unrelated labels rather than neutral labels. To evaluate the bias scores for the LLM $M$ under investigation, we calculate the bias score $\mu_{d,M}$ for social dimension $d$ where $d\in$race, gender, religion, profession, unrelated} as
$\mu_{d,M}=\frac{1}{|\mathcal{P}_{M}|}\sum_{p\in\mathcal{P}_{M}}\max_{s\in p}(\mu_{d,s})$
where $\mathcal{P}_{M}$ is the set of passages generated from LLM $M$ using the prompt-library, $p$ is a passage in $\mathcal{P}_{M}$, $s$ is a sentence in $p$ and $\mu_{d,s}$ is the bias score given to each sentence. The bias score is the probability of stereotype bias predicted by the proposed sentence-level stereotype detector for each social dimension. We further employ a deviation-based normalisation since we have observed a consistently increase in the unrelated score from GPT-2 to GPT-3.5 to GPT-4. To provide a neutral view of an LLM’s bias, we analyze their deviation from unrelated scores across all types of stereotypes. The deviation $\Delta_{d,M}$, for each social dimension $d$, defined by the equation:
$\Delta_{d,M}=\mu_{d=\text{unrelated},M}-\mu_{d\neq\text{unrelated},M}$
This deviation $\Delta_{d,M}$ quantifies the difference between the stereotype bias score in the ”unrelated” category and the scores in other categories for the same model. A significantly low negative value of $\Delta_{d,M}$ suggests a lower presence of bias (i.e., the bias score is normally less than the unrelated score), indicating a more unbiased output. Conversely, a deviation that is closer to zero or less negative points to a bias score that closely approximates the unrelated baseline, signifying a greater degree of bias.
@@ 4 Results and Discussion
The superior performance of multi-class models underscores the significance of stereotype intersectionality-a concept introduced by (Crenshaw, 1989), highlighting the critical need to consider multiple intersecting stereotypes simultaneously. Our finding affirms this, showing that understanding one stereotype dimension improves the detection of others. This finding is also supported by (Ali et al., 2020), which demonstrated that Multi-Task Learning (MTL) methods, when carefully fine-tuned, outperform the Single-Task Learning (STL)s methods. Nonetheless, our approach faces challenges, particularly in model generalizability across evolving stereotypes, a complexity highlighted by (Yoon, 2020). This suggests that, while promising, multi-dimensional training must evolve to keep pace with the dynamic nature of stereotypes, marking a crucial direction for future research.
@@
Stereotype Type Training Setting Macro Precision Macro Recall Macro F1 Score
@@
Testing Data Training Data Performance Metrics
Macro F1 Score Weighted F1 Score Accuracy
StereoSet StereoSet 0.744 0.807 0.808
CrowsPairs 0.165 0.188 0.411
MGSD 0.747 0.812 0.812
CrowsPairs StereoSet 0.184 0.293 0.599
CrowsPairs 0.568 0.845 0.864
MGSD 0.543 0.726 0.789
MGSD StereoSet 0.728 0.797 0.797
CrowsPairs 0.177 0.202 0.426
MGSD 0.743 0.809 0.809
@@
Baseline Method Macro Precision Macro Recall Macro F1 Score Complexity
Random Selection 0.112 0.114 0.091 -
ZeroShot Classification 0.394 0.372 0.314 184M
LogReg (TF-IDF) 0.514 0.472 0.494 Dependent
KSVM (TF-IDF) 0.533 0.481 0.505 Dependent
@@
To address RQ4, we present explainability visualizations in Figures 1, 2, and 3, which illustrate the decision-making process of our BERT-based classifiers in specific to ALBERT-V2. Using the example sentence ”The baby loved the presence of the caring mommy,” we explore the model’s interpretation. The SHAP visualization in Figure 1 suggests that the phrase ”the caring mommy” significantly influences the model’s prediction, indicating a stereotype linked to gender. The word ”caring” also stands out, reinforcing the stereotype associated with nurturing roles. In Figure 2, LIME quantifies the impact of individual words, with ”caring” and ”mommy” showing the most substantial positive influence on the stereotypical prediction. The word ”loved” has a smaller effect, likely due to its association with ”caring” and ”mommy”. BERTViz, depicted in Figure 3, provides a detailed look at the model’s attention mechanisms. Specifically, Layer 5, Head 5 displays focused attention between ”caring” and ”mommy,” indicating that these words’ stereotypical relevance captures the model’s attention. This pattern persists across multiple layers and heads, offering insight into how BERT prioritizes information associated with gender stereotypes.
@@
The analysis reveals that SHAP, LIME, and BERTViz are in agreement and align with our human understanding of gender stereotypes, validating our model’s effectiveness in identifying stereotype-indicative words like ”caring” and ”mommy.” Additional examples can be found in the Appendix A.1. However, the reliance on theses tools for model validation presents challenges. Specifically, SHAP and LIME offer local approximations of the model’s behavior, potentially not reflecting its global decision-making process accurately. These methods’ effectiveness heavily relies on the correct selection and representation of features. Misleading explanations may result if the model relies on abstract patterns that are challenging to interpret in human terms. Moreover, there is an ongoing debate about the explainability of attention mechanisms, such as those visualized by BERTViz, in providing meaningful explanations for model predictions. Some research suggests that attention might not always correlate with feature importance (Jain & Wallace, 2019).
To answer RQ5, Tables 4 reveals some key findings. The Table 4 presents their average deviations from unrelated scores for models in the GPT family along with deviation and bias scores for each stereotype dimensions. First, no single model excels in every category, highlighting the complexity and trade-off of completely eliminating bias. However, there is a clear trend: as we move from GPT-2 to GPT-4, the average deviation generally decreases. This is most evident in the ’Race’ category, where the deviation dropped from -0.018 in GPT-2 to -0.105 in GPT-4. The Average Deviation shows a consistent decline across model generations. These trends indicate that while no model is perfect, debiasing in LLMs are making them less biased over time. While progress has been observed in mitigating race and gender stereotypes, biases related to profession and religion have not seen comparable advancements. This means the debiasing is not uniformly distributed across all types of stereotypes. We hypothesize that this discrepancy may be due to a more focused attention on race and gender, possibly at the expense of adequately debiasing in profession and religion, suggesting an oversight in certain dimensions.
@@
5 Conclusion and Future Work
In conclusion, we have developed a foundation for a framework for auditing bias in LLMs through text-based stereotype classification. Using the MGS Dataset and fine-tuned PLMs, our approach surpasses proposed baselines and demonstrates the superiority of multiple stereotype dimensions classifiers over the single stereotype dimension ones and the superiority of use MGS Dataset over the single stereotype dataset. To verify the decisions made by our models, we incorporated XAI techniques such as SHAP, LIME, and BertViz. Benchmark results further confirm a reduction in bias in newer versions of the GPT series.
For future work, first, we aim to discover multi-label dataset and model development to detect overlapping stereotypes and evaluate their synergistic effect on efficacy, moving beyond the the current multi-class approach. Secondly, we plan to expand stereotype categories to include for example, LGBTQ+ (WinoQueer (Felkner et al., 2023)) and regional stereotypes (SeaGULL Jha et al. (2023)). Additionally, we will discover token-level stereotype detection to improve analysis granularity, inspired by token-level hallucination detection (Liu et al., 2021a).
Ethics Statement

### v2

# Stereotype Detection in LLMs: A Multiclass, Explainable, and Benchmark-Driven Approach

@@ Abstract
Stereotype detection is a challenging and subjective task, as certain statements—such as "Black people like to play basketball"—may not appear overtly toxic but still reinforce racial stereotypes. With the increasing prevalence of Large Language Models (LLMs) in human-facing Artificial Intelligence (AI) applications, detecting such type of biases is essential. However, LLMs risk perpetuating and even exacerbating stereotypical outputs derived from their training data. A reliable stereotype detector is crucial not only for benchmarking bias but also for monitoring model I/O, filtering training data, and ensuring fairer model behavior in downstream applications.
This paper first introduces the Multi-Grain Stereotype (MGS) dataset, consisting of 51,867 instances across gender, race, profession, religion, and other stereotypes, curated from multiple existing datasets. We evaluate various machine learning approaches to establish baselines and fine-tune language models of different architectures and sizes, presenting a suite of stereotype multiclass classifiers trained on the MGS dataset. Given the subjectivity of stereotype, explainability is essential to align model learning with human understanding of stereotypes. We employ eXplainable AI (XAI) tools, including SHAP, LIME, and BertViz, to assess whether the model’s learned patterns align with human intuitions about stereotype. Additionally, we develop stereotype elicitation prompts and benchmarking the presence of stereotypes in text generation tasks using popular LLMs, employing best-performing stereotype classifiers.
Overall, our experiments reveal key findings: i) Multi-dimensional classifiers outperform single-dimension classifiers, ii) the integrated MGS dataset significantly improves both in-dataset and cross-dataset generalization compared to individual datasets, and iii) Newer versions of GPT-family LLMs demonstrate reduced stereotypical content generation; however, no model excels in all categories, highlighting the challenge of fully eliminating bias across different dimensions.  [^1] The code will be made available upon acceptance.
Stereotype Detection in LLMs: A Multiclass, Explainable, and Benchmark-Driven Approach
Zekun Wu1,2, Sahan Bulathwela2, María Pérez-Ortiz2, Adriano Koshiyama1 1Holistic AI, 2University College London Correspondence: zekun.wu@holisticai.com, adriano.koshiyama@holisticai.com
@@ 1 Introduction
The field of Artificial Intelligence (AI) continues to evolve with Large Language Models (LLMs) showing both potential and pitfalls. This research explores the ethical dimensions of LLM auditing in Natural Language Processing (NLP), with a focus on text-based stereotype classification and bias benchmarking in LLMs. The advent of state-of-the-art LLMs including OpenAI’s GPT series (Brown et al., 2020; Radford et al., 2019; OpenAI, 2023), Meta’s LLaMA series (Touvron et al., 2023a, b), and the Falcon series (Almazrouei et al., 2023) has magnified the societal implications. These LLMs, shown up with abilities like in-context learning as a few-shot learner (Brown et al., 2020), reveal emergent capabilities with increasing parameter and training token sizes (Wei et al., 2022). However, they show fairness concerns due to their training on extensive, unfiltered datasets such as book (Zhu et al., 2015) and Wikipedia corpora (Foundation, 2021b), and large internet corpora like Common Crawl (Foundation, 2021a). This training data often exhibits systemic biases and could further lead to detrimental real-world effects, confirmed by studies (May et al., 2019a; Bordia and Bowman, 2019; Davidson et al., 2019; Magee et al., 2021). For instance, biases in LLMs and AI systems can reinforce political polarization as seen in Meta’s news feed algorithm (Bakshy et al., 2015), and exacerbate racial bias in legal systems as documented in predictive policing recidivism algorithms like COMPAS (Angwin et al., 2016). Furthermore, issues such as gender stereotyping and cultural insensitivity are highlighted by tools like Google Translate and Microsoft’s Tay (Prates et al., 2019; Neff and Nagy, 2016). Most existing studies focus on either bias benchmarks in LLMs or text-based stereotypes detection and overlook the interaction between them, which remains underexplored and indicates gaps. Our study makes a clear line between Bias, as observable deviations from neutrality in LLM downstream tasks, and Stereotype, a subset of bias entailing generalized assumptions about certain groups in LLM outputs. Aligning with established stereotype benchmark: StereoSet (Nadeem et al., 2020), we detect text-based stereotypes at sentence granularity, across four societal dimensions—Race, Profession, Religion, and Gender—within text generation task conducted with LLMs.
@@ 2 Related Works
Further expanding on the benchmark, StereoSet (Nadeem et al., 2020) and CrowS-Pairs (Nangia et al., 2020) are popular dataset-based stereotype benchmarking approaches that use the examples in the datasets to calculate the masked token probabilities and pseudo-likelihood-based scoring of the LLM to assess whether stereotypical results are output. A key disadvantage of these approaches is that the stereotype assessment’s generalisation bounds are limited to the diversity of the examples in the datasets. On the contrary, we use these examples to fine-tuned models to detect stereotypes from any generated text. This gives our approach the advantage of assessing the LLM’s bias based on any text output generated by the LLM rather than within the constraints of the labelled datasets. Benchmarks such as WinoQueer (Felkner et al., 2023) and SeeGULL (Jha et al., 2023) focus on stereotype types that are out of the scope of this work (e.g. regional stereotype etc.). Benchmarks such as WEAT (Caliskan et al., 2017) and SEAT (May et al., 2019b) use pre-defined attribute and target word sets to assess stereotypical language, making them similar to StereoSet (Nadeem et al., 2020), WinoQueer (Felkner et al., 2023), and CrowS-Pairs (Nangia et al., 2020) approaches exposed to the same limitations. Meanwhile, the BBQ benchmark (Parrish et al., 2022), which centers on a question-answering task, has been observed to prompt LLMs to avoid answering due to its explicit question design. BOLD (Dhamala et al., 2021), focusing on text generation, employs various metrics but lacks a solid measure to identify stereotypes. These prior collectively reveal a significant gap in current methods, as they do not facilitate a high granular stereotype detection in freely generated text by LLMs.
Additionally, several other prior works (Bolukbasi et al., 2016; May et al., 2019a) could be used to implement token-level stereotype detection that is out of scope for this work as we focus on sentence-level stereotype detection. Albeit, these works also lack transparency, a gap our work addresses through eXlainable AI (XAI) techniques. While emerging LLM evaluation frameworks like DeepEval (Team, 2023), HELM (Liang et al., 2022), and LangKit (whylabs, 2023) takes a holistic view on bias evaluation, our framework complements them as our proposal can become a subcomponent within their systems.
@@ 3 Methodology
  • RQ3: How does our multiclass classifier built for stereotype detection compare to other relevant baselines?
@@ 3.1 MGS Dataset (RQ1-3)
The Multi-Grain Stereotype Dataset (MGSD) was constructed from two crowdsourced sources, StereoSet Nadeem et al. (2020) and CrowS-Pairs Nangia et al. (2020). It consists of 51,867 instances, divided into training and testing sets in an 80:20 ratio. This division ensures stratified sampling across different stereotype types and data sources, providing a diverse range of examples for model creation and facilitating multiclass learning. The dataset includes columns for original text (text), text with marked stereotypical words/phrases (text_with_marker), stereotype type (stereotype_type), stereotype level (category), original data source (data_source), and gold label (label).
The labelling scheme in the MGS Dataset classifies texts into three categories: "stereotype", "neutral", and "unrelated", covering four social stereotype dimensions: "race", "religion", "profession", and "gender". There are nine label types: "unrelated", "stereotype_race", "stereotype_gender", "stereotype_profession", "stereotype_religion", "neutral_race", "neutral_gender", "neutral_profession","neutral_religion". The dataset supports both sentence-level and token-level classification tasks. For preprocessing, the text was tokenized, and "$=$" markers were inserted to highlight stereotypical tokens (e.g., He is a doctor $\rightarrow$ He is a $=$doctor$=$). This approach allows for the future use of the dataset in training token-level stereotype detectors and aids in generating prompts and counterfactual scenarios for sentence-level detector model evaluations.
StereoSet includes two types of data examples: (i) intra-sentence, where bias is within a single sentence, and (ii) inter-sentence, where bias spreads across multiple sentences. (iii) CrowS-Pairs features pairs of sentences that carry the stereotype or stereotype with counterfactual identity/group. In (i) intra-sentence cases, the correlated label is assigned to the single sentence. In (ii) inter-sentence cases, sentences are merged, and the label is assigned to create the final MGS dataset. For (iii) CrowS-Pairs, the first pair of sentences and the "sent_more" feature are chosen, and the label is assigned to finalize the dataset. The MGS Dataset serves as a valuable resource for evaluating and understanding stereotypes. We conducted exploratory data analysis and provided details in Appendix A.4.
@@ 3.2 Finetuning the Stereotype Classifier and Explaining It (RQ1-4)
For RQ3, in order to compare the new model with comparative baselines, we developed popular machine learning methods as baselines, in addition to the range of PLMs we had fine-tuned. This was necessary as we could not find suitable multiclass baselines in previous research. We implemented the i) Random selection, that assigns labels at random, ii) a Logistic regression, and iii) Kernel SVM (sigmoid kernel identified empirically) models trained TF-IDF features from MGS Dataset. iiii) Finally, zero-shot classification task, we choose the "sileod/deberta-v3-base-tasksource-nli" DeBERTa-based model that has shown the best performance in this task(Sileo, 2023).
For RQ4, to ensure robust validation and interpretation of our stereotype classifier, we employ multiple explainability ai methods for feature attribution and model structural interpretability. This allows us to check for consistency of explanations as different explainability methods can yield varying results in feature importance (Swamy et al., 2022). Specifically, we apply SHAP (Lundberg and Lee, 2017) and LIME (Tulio et al., 2016), two popular model-agnostic explainability techniques, to identify the text tokens most influential in the classification process. We use randomly selected examples from the test set of MGS Dataset to analyse explanations. Additionally, we utilize BERTViz (Vig, 2019), a model-specific visualization tool for transformer models, to observe how the model’s attention heads engages with specific tokens across layers.
@@ 3.3 Stereotype Elicitation Experiment and Bias Benchmarks (RQ5)
For RQ5, we developed a method for benchmark prompt creation, resulting in a prompt library in Appendix A.2, which designed for effectively elicits stereotypical text from the LLMs. We take data examples from the MGS dataset and use the token level text markers to identify the prompts (the part of the example before the marker) for the LLM under investigation. When selecting examples for generating prompts, we use word count-based prioritization logic, where initially, we target long examples resulting in detailed prompts. We generate prompts from the dataset for different societal dimensions ($\approx 200$ per dimension). We further validate the neutrality of the identified prompts using the proposed model to ensure that all prompts have been classified as "unrelated." Then, we use the prompt library to probe the LLM under investigation to complete the rest of the passage (prompt). To evaluate the effectiveness of our prompts in eliciting stereotype-related content, we conducted perplexity tests, as documented in Appendix A.3. These tests compare the LLMs’ responses against a stratified sample from the MGS Dataset, based on labels. Finally, we use the generated output to detect stereotypes, which forms the final assessment.
In this paper, we assess bias in the GPT series of LLMs, focusing on stereotype and unrelated labels. To compute the bias score $\mu_{d,M}$ for social dimension $d\in$race, gender, religion, profession, unrelated}, we use:
$\mu_{d,M}=\frac{1}{|\mathcal{P}_{M}|}\sum_{p\in\mathcal{P}_{M}}\max_{s\in p}(\mu_{d,s}),$
where $\mathcal{P}_{M}$ represents the set of passages generated from LLM $M$, and $\mu_{d,s}$ is the stereotype bias probability assigned by the detector for each sentence.
Given the rise in unrelated scores from GPT-2 to GPT-4, we normalize the bias score using a deviation metric to compare stereotype bias against unrelated content. The deviation $\Delta_{d,M}$ for each social dimension $d$ is:
$\Delta_{d,M}=\mu_{d=\text{unrelated},M}-\mu_{d\neq\text{unrelated},M},$
where more negative values of $\Delta_{d,M}$ indicate lower bias, and values closer to zero suggest greater bias. This normalization accounts for model tendencies toward neutrality, making bias comparisons across models more meaningful.
@@ 4 Results and Discussion
The superior performance of multiclass models underscores the significance of stereotype intersectionality-a concept introduced by (Crenshaw, 1989), highlighting the critical need to consider multiple intersecting stereotypes simultaneously. Our finding affirms this, showing that understanding one stereotype dimension improves the detection of others. This finding is also supported by (Ali et al., 2020), which demonstrated that Multi-Task Learning (MTL) methods, when carefully fine-tuned, outperform the Single-Task Learning (STL)s methods. Nonetheless, our approach faces challenges, particularly in model generalizability across evolving stereotypes, a complexity highlighted by (Yoon, 2020). This suggests that, while promising, multi-dimensional training must evolve to keep pace with the dynamic nature of stereotypes, marking a crucial direction for future research.
@@
Stereotype Setting Precision Recall F1
@@
Test Data Train Data Performance Metrics
F1 Accuracy
StereoSet StereoSet 0.744 0.808
CrowsPairs 0.165 0.411
MGSD 0.747 0.812
CrowsPairs StereoSet 0.184 0.599
CrowsPairs 0.568 0.864
MGSD 0.543 0.789
MGSD StereoSet 0.728 0.797
CrowsPairs 0.177 0.426
MGSD 0.743 0.809
@@
Method Precision Recall F1 Size
Random 0.112 0.114 0.091 -
Zero-Shot 0.394 0.372 0.314 184M
LogReg TF-IDF 0.514 0.472 0.494 -
KSVM TF-IDF 0.533 0.481 0.505 -
@@
To address RQ4, we present explainability visualizations in Figures 1, 2, and 3, which illustrate the decision-making process of our BERT-based classifiers in specific to ALBERT-V2. Using the example sentence "The baby loved the presence of the caring mommy," we explore the model’s interpretation.
@@
The SHAP visualization in Figure 1 shows that the phrase "the caring mommy" strongly influences the model’s prediction, highlighting a gender-linked stereotype. The word "caring" stands out, reinforcing the association with nurturing roles. Similarly, in Figure 2, LIME quantifies word impacts, with "caring" and "mommy" having the highest positive influence on the stereotypical prediction. The word "loved" has a smaller effect, likely due to its link with "caring" and "mommy." BERTViz, in Figure 3, reveals that Layer 5, Head 5 focuses attention between "caring" and "mommy," showing how the model emphasizes words related to gender stereotypes across multiple layers and heads.
@@
The analysis reveals that SHAP, LIME, and BERTViz are in agreement and align with our human understanding of gender stereotypes, validating our model’s effectiveness in identifying stereotype-indicative words like "caring" and "mommy." Additional examples can be found in the Appendix A.1. However, the reliance on theses tools for model validation presents challenges. Specifically, SHAP and LIME offer local approximations of the model’s behavior, potentially not reflecting its global decision-making process accurately. These methods’ effectiveness heavily relies on the correct selection and representation of features. Misleading explanations may result if the model relies on abstract patterns that are challenging to interpret in human terms. Moreover, there is an ongoing debate about the explainability of attention mechanisms, such as those visualized by BERTViz, in providing meaningful explanations for model predictions. Some research suggests that attention might not always correlate with feature importance (Jain and Wallace, 2019).
To answer RQ5, Table 4 reveals key trends in bias reduction across GPT models. It presents the average deviations from unrelated scores for each model, along with bias and deviation scores for race, gender, profession, and religion. No model excels in all categories, highlighting the challenge of fully eliminating bias across dimensions. However, a clear trend of improvement is evident from GPT-2 to GPT-4, particularly in the ’Race’ dimension, where the deviation dropped from -0.018 in GPT-2 to -0.105 in GPT-4. GPT-4 also shows the lowest overall average deviation at -0.173, indicating it is progressively less biased than its predecessors. The average deviations across all stereotype dimensions consistently decrease as models advance, demonstrating that debiasing efforts have been effective, though unevenly distributed. Significant progress has been made in mitigating race and gender biases, but biases in profession and religion have not improved to the same extent. For example, while GPT-4 shows substantial reductions in race and gender biases, deviations for profession and religion remain closer to their unrelated baselines. We hypothesize this imbalance may result from a focused effort on race and gender biases, potentially at the expense of profession and religion. This suggests a need for a more comprehensive, evenly distributed debiasing strategy. Additionally, the unrelated score consistently rises from GPT-2 to GPT-4, with GPT-4 showing the highest unrelated score of 0.834, indicating a greater tendency to produce neutral text. The comprehensive deviation score metric normalizes bias scores against this tendency, providing a clearer view of the progress in debiasing LLMs.
5 Limitation and Future Work
Despite the contributions of our framework, several limitations exist. First, our deviation-based bias metric relies on the unrelated score as a baseline, which may not capture all forms of bias. Vocabulary discrepancies between the prompt library and classifier training data could impact stereotype detection accuracy. Although we mitigated this through vocabulary checks and human evaluation, further validation is required to ensure robustness across different linguistic contexts. While BertViz offers useful insights, it has limitations in providing reliable explanations. However, we observed alignment between patterns identified by BertViz, LIME, and SHAP, reinforcing interpretability consistency. Future work should explore alternative methods for more reliable explanations, particularly in complex cases where attention-based tools may fall short. Additionally, the model used for stereotype classification has not been fully calibrated, meaning the scores should not be interpreted as true probabilities. Future work must address this by properly calibrating the model to improve the reliability of bias quantification. The uneven bias mitigation across dimensions, with more progress in race and gender than profession and religion, requires further investigation to understand whether this arises from dataset distribution or model architecture biases.
Future directions include developing multi-label models to detect overlapping stereotypes, expanding stereotype categories to underrepresented groups like LGBTQ+ (WinoQueer (Felkner et al., 2023)) and regional biases (SeaGULL (Jha et al., 2023)), and exploring token-level stereotype detection inspired by token-level hallucination detection (Liu et al., 2021a). Additionally, more rigorous human evaluation, incorporating diverse annotator pools, will be essential for improving the reliability of future bias audits.
6 Ethics Statement